OpenStudy (anonymous):

Find a power series representation centered at (a) for the function and determine all values of x where the power series converge

OpenStudy (anonymous):

\[\frac{ (x+2)^2 }{ 5-x} ,a=-2\]

OpenStudy (anonymous):

I set t=x+2 x=t-2 ((t-2)+2)^2/(5-(t-2) t^2/-t+7

OpenStudy (anonymous):

IDK what to do after this

OpenStudy (anonymous):

\[-\frac{ t }{ 7 }\]???